OpenStudy (anonymous):

sqrt{36-4x ^{2}} what is the domain and how do you write it in interval?

OpenStudy (anonymous):

\[\sqrt{36-4x^2}\] right? this means \[36-4x^2\geq0\] \[9-x^2\geq 0\] now \[y=9-x^2\] is a parabola that faces down, so it will be negative between the zeros and positive outside them. your answer is \[-3\leq x\leq 3\] or \[[-3,3]\]

OpenStudy (anonymous):

Thanks Satalite73!!